Prev Next

SW&IT Meeting Minutes 2020-03-09

Attendees: David, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200224
LR20200224 - Bec, Matthieu 80.0
LR20200224 - Holmberg, Carl 40.0
LR20200224 - Edmeades, David 80.0
LR20200224 - Sola, Igor 40.0
LR20200224 - Hooper, Stephen 80.0
LR20200224 - Zhang, Yang 72.0
LR20200224 - Zhang, Xianyu 79.0

News from management

  • COVID-19 and travel limitations
  • new office plan layout, return to 5th floor

Previous Action Items

Owner Action Due Date Completion Date
DE timecapsule obsdev 2020-02-12  
DE oVirt for PEPSI 2020-02-12  

Open topics

  • DX hexapod for ASM tbc. Wednesday
  • state of the aluminizing PC?
  • UITS reloc?

Team reports

YZ

  • Updated LBC tech chip move script (will create github pull request).
  • Completed a draft MODS data cleaning script (IT#8061), tested it with Olga (works well, have a few enhancements to make, a question: will a permanent NFS mount on MODS data machine cause stability issues? since MODS data machines are critical MODS equipments).
  • AO-Cravat meeting with Al Conrad (discussed about automated report generation, currently ELab IDL code is the hardest part).
  • Safety training.
  • Next to do:
    • Review Al's existing AO-Cravat OSCO data processing material and prepare for a next discussion.
    • Organize existing C and Python code for Strehl calculation.
    • Follow up with Olga regarding MODS data cleaning.
    • Work with Olga on MODS 64-bit programs issues.
    • Look at LBC IT#7854 (Possibility of turning off HK via UI while all other systems are still ON).
    • Get a trial Mathematica for TMS stuffs.

DE

  • Summit laptop almost ready to go up
    • User setup for Wiki; do we need any restricted pages to be available?
  • Veeam steady state
  • VMHost1 array rebuilt
  • VMHost2 array rebuild this week, pending receipt of new SSD
  • AB laptop arrived
  • Mac Mini delivery delayed; CoViD-19 related?
  • Nasty malware vector masquerading as a certificate problem, Windows/OSX payload: https://www.zdnet.com/article/backdoor-malware-is-being-spread-through-fake-security-certificate-alerts/
  • VMHost1 reboot to clear consolidation problems/upgrade hypervisor after 2's array rebuilt
  • USB-3 fiber extender optics module not changeable; gave thread to LeRoy.

MB

  • Last week
    • OC, WPC, management
    • board interview
    • dms (ovms psd, asm, refmem, misc.)
    • Shark-VIS AT review
  • This week
    • rowg#3
    • Remote Ops and current pandemic, particularly affecting our Italians partners
    • swgdev01 ks f30
    • dms ASM with XZ

CH

  • Last week:
    • AHS: Added write to EPICS for state change to LBTI AHS client.
  • This week:
    • Vacation

SH

  • toyed with DMS to replace wda.sh
  • SSL certs for wiki.lbto.org, VPNs
  • replaced video card in 507-3, switched 507-1 to using X
  • obs4 issues
  • ssh keys
  • fiber meeting
  • discussion with DR about machine for Ignition

next:

  • moving back to five
  • more DMS and info stuff
  • look a little more into Xpra
  • upgrade DNS
Topic revision: r10 - 09 Mar 2020, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback